# Haoyu Zhang12  Qiaohui Chu12  Meng Liu3  Yunxiao Wang4  Bin Wen5  Fan Yang5 Tingting Gao5  Di Zhang5  Yaowei Wang12  Liqiang Nie1

### v1

@@ 4.4 Stage 3: Student Self-Practice
In the final stage, we focus on enhancing the instruction-following capabilities of MLLMs to address diverse egocentric downstream tasks. To achieve this, we employ the instruction-tuning data EgoIT, as described in Section 3.2. By concatenating the representations of the egocentric sample $x$ and its corresponding mapped exocentric sample $F(x)$, we create a comprehensive semantic input for the LLM. To ensure the LLM aligns its responses more effectively with task instructions, we apply Low-Rank Adaptation (LoRA) [28] to the frozen LLM, tuning it alongside the visual encoder and the mapping function $F$ using VTG loss.
5 Experiment
Source Videos Clips Sec Questions
EgoSchema 5,031 5,031 180.0 5,031
QAEgo4D 166 1,854 310.0 1,854
[…]
EgoPlan Val 678 3,339 668.0 3,339
VLN-QA 200 200 7.5 200
[…]
Table 2: Overview of the evaluation sources in EgoBench. Tasks requiring additional processing are marked with $\ddagger$.
Method ViEncoder LLM Frame Eval EgoSchema QAEgo4D EgoTaskQA
Acc. Acc. Acc./Score Acc. Acc.
VLP Models
EgoVLP [49] TS-B - 16 FT - 22.6$\dagger$ 17.3/1.9$\dagger$ 42.5 38.7
EgoVLPv2 [63] TS-B - 16 FT - 25.8$\dagger$ 19.7/2.1$\dagger$ 46.3 42.3
MFAS [95] TS-B - 16/32 FT - 37.1$\dagger$ 23.2/2.2$\dagger$ 48.7 45.4
GroundVQA [18] TS-B+IV - 16 FT - 50.2 29.0/2.6$\dagger$ - -
Closed-source MLLMs
Gemini 1.0 Pro [76] - - - ZS 55.7 - - - -
GPT-4V [61] - GPT-4 - ZS 55.6 - - - -
Open-source MLLMs
LLaMA-VID [46] EVA-G Vicuna-7B 1fps ZS 38.5 38.6$\dagger$ 23.5/2.0$\dagger$ 29.0$\dagger$ 31.6$\dagger$
Video-LLaVA [48] LB Vicuna-7B 8 ZS 38.4 35.2$\dagger$ 28.7/2.7$\dagger$ 29.8$\dagger$ 30.6$\dagger$
VideoChat2 [43] UMT-L Vicuna-7B 16 ZS 42.2 43.9$\dagger$ 15.9/1.3$\dagger$ 33.9$\dagger$ 37.3$\dagger$
VideoLLaMA2 [11] CLIP-L-336 Mistral-7B 8 ZS 50.5 47.7$\dagger$ 26.7/2.6$\dagger$ 40.4$\dagger$ 42.3$\dagger$
VideoLLaMA2 [11] CLIP-L-336 Mistral-7B 16 ZS 51.7 48.5$\dagger$ 25.3/2.6$\dagger$ 42.1$\dagger$ 42.6$\dagger$
VideoChat2 [43] IV Mistral-7B 16 ZS 54.4 45.4$\dagger$ 21.9/1.7$\dagger$ 42.3$\dagger$ 47.4$\dagger$
Exo2Ego CLIP-L-336 Mistral-7B 16 ZS 61.3 62.1 28.3/2.7 44.7 50.3
Table 3: Performance comparison with the latest methods on the EgoSchema, QAEgo4D, and EgoTaskQA datasets. “TS”, “IV”, and “LB” refer to TimeSformer, InternVideo, and LanguageBind, respectively. “FT” and “ZS” indicate fine-tuning and zero-shot settings. The QAEgo4D dataset is categorized into close and open question answering, while the EgoTaskQA dataset includes direct and indirect subsets. The best results among open-source MLLMs are highlighted in bold, and $\dagger$ denotes inference results via open-source code.
Method Charades-Ego EK-100 MIR
mAP mAP nDCG
EgoVLP [49] 25.0 26.0 28.8
LAVILA-L [99] 28.9 40.0 36.1
EgoVLPv2 [63] 26.2 - -
LLaMA-VID [46] 31.0$\dagger$ 46.1$\dagger$ 59.3$\dagger$
Video-LLaVA [48] 43.5$\dagger$ 45.2$\dagger$ 58.6$\dagger$
VideoChat2 [43] 57.0$\dagger$ 42.6$\dagger$ -
VideoLLaMA2 [11] 68.1$\dagger$ 47.8$\dagger$ 60.6$\dagger$
Exo2Ego 70.9 49.7 63.6
[…]
Method EgoPlan Val VLN-QA
Acc. Acc.
Video-ChatGPT [54] 27.3 29.5
VideoChat [42] 27.5 23.5
mPLUG-Owl [90] 27.8 26.0
Otter [39] 28.1 23.5
Video-LLaMA [93] 28.6 30.0
GPT-4V [61] 38.0 31.0
LLaMA-VID [46] 27.2$\dagger$ 32.3$\dagger$
Video-LLaVA [48] 27.5$\dagger$ 34.0$\dagger$
VideoChat2 [43] 26.1$\dagger$ 35.0
VideoLLaMA2 [11] 29.3$\dagger$ 37.5$\dagger$
Exo2Ego 42.7 44.5
Table 5: Zero-shot performance comparison with the latest methods on the EgoPlan and VLN-QA datasets. The best results are highlighted in bold, and $\dagger$ indicates inference results obtained through open-source code.
5.1 Implementation Details
@@
Datasets. The datasets employed in each stage of training are summarized in Table 1, showcasing a substantial amount of vision-text data across various types. After training, we evaluated our model using the proposed EgoBench benchmark, which includes eight tasks with multiple configurations: EgoSchema [58] for reasoning, QAEgo4D [6] for episodic memory, Charades-Ego [75] for action recognition, EPIC-KITCHENS-100 [15] for multi-instance retrieval, VLN-QA [38] for egocentric navigation, EgoPlan [10] for planning, EgoMCQ [49] for alignment, and EgoTaskQA [31] for hybrid tasks. Detailed descriptions of these evaluation sources are provided in Table 2.
@@
Model. Our model is built upon the VideoLLaMA2 [11] baseline. We consistently utilized the CLIP-Large-336555https://huggingface.co/openai/clip-vit-large-patch14-336. as the visual encoders and Mistral-7B-Instruct [^6] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2. as the LLM. The mapping functions $F$ and $G$ consist of nine ResNet [26] blocks between the down-sampling and up-sampling operations, sourced from [102]. In Stage 3, we incorporated LoRA [27] into the LLM, with a rank set to 128, an alpha value of 256, and a dropout rate of 0.1. All experiments are conducted using the PyTorch [62] framework on 16 A800 GPUs. Further experimental details, including benchmark construction and hyperparameter settings, are provided in Supp. C and D.1.
5.2 Quantitative Evaluation
Method EgoMCQ
Acc. (Inter) Acc. (Intra)
LLaMA-VID [46] 56.5$\dagger$ 23.5$\dagger$
Video-LLaVA [48] 60.9$\dagger$ 25.1$\dagger$
HelpingHands [91] 80.2$\dagger$ 32.5$\dagger$
AVION [98] 81.8$\dagger$ 33.6$\dagger$
VideoLLaMA2 [11] 82.1$\dagger$ 32.0$\dagger$
VideoChat2 [43] 85.1$\dagger$ 37.6$\dagger$
Exo2Ego 88.4 41.2
Table 6: Zero-shot performance comparison with the latest methods on the EgoMCQ dataset. “Inter” refers to cases where the candidates and input clips come from different videos, while “intra” indicates that they are from the same video, presenting a more challenging scenario.
@@ 5.2 Quantitative Evaluation
Table 7: Performance comparison of different model architecture. The LLM is fixed at Mistral-7B-Instruct-v0.2, the symbol “✰” denotes replacing the existing modules with fully connected networks, and “Avg” represents the average performance across all metrics in Figure 1.
Tables 3, 4, 5, and 6 present the experimental results across various sources within the EgoBench benchmark, which can be analyzed as follows:
  • Table 3 lists evaluation results across various tasks, including reasoning and episodic memory. These results indicate that MLLMs, in a zero-shot setting, achieve performance comparable to Vision-Language Pre-training (VLP) models, highlighting their strong generalisability across data domains. However, they underperform in the egocentric domain, as evidenced by VideoChat2 achieving 70.0% on MSVD-QA [87], compared to only 21.9% on QAEgo4D. Our Exo2Ego method significantly surpasses all open-source MLLMs and GPT-4V under almost all settings across three egocentric tasks.
  • Table 4 presents zero-shot results for action recognition on Charades-Ego and multi-instance retrieval on EPIC-KITCHENS-100. Our Exo2Ego method demonstrates significant improvements over VLP methods and open-source MLLMs across all metrics for both tasks, validating its effectiveness in understanding the details of hand-object interactions.
  • Planning and navigation are critical tasks in egocentric video understanding domain. Table 5 summarizes our method’s performance against existing MLLMs. The results reveal that open-source MLLMs exhibit accuracy rates concentrated between 26.0% and 29.0% on EgoPlan, highlighting their difficulties for planning. In contrast, our method significantly outperforms all competitors, including GPT-4V (4.7% and 13.5% absolute gains), demonstrating that Exo2Ego framework can effectively tackle diverse egocentric tasks.
  • Video-text matching, a fundamental alignment task, is also essential for evaluation. As the results presented in Table 6, our Exo2Ego approach consistently achieves optimal performance across different settings, further validating its superiority for visual descriptions.
These comparisons collectively highlight the effectiveness and generalizability of our Exo2Ego method across a wide range of egocentric video understanding tasks.
5.3 Ablation Study
In this section, we conducted a comprehensive analysis of how our model architecture, parameter update, training data, and training strategy impact overall performance. More experimental results are available in Supp. D.2.
Model Architecture. As summarized in Table 7, we investigated the impact of the mapping functions, LoRA, and loss functions on model performance. Comparing results 1 and 2, we observed that the introduction of LoRA significantly improves instruction-following capabilities, leading to enhanced performance across various tasks. This highlights LoRA’s effectiveness in refining the model’s ability to follow complex instructions. In the comparison between results 1 and 3, we found that the mapping $G$ and CCL loss play a critical role in establishing an effective mapping between the egocentric and exocentric domains, thereby improving performance on downstream egocentric tasks. In the fourth experiment, we removed the KL loss, which incorporates guidance from exocentric data. The performance gap between this result and result 1 reinforces the importance of exocentric knowledge in enhancing egocentric video understanding. In the fifth experiment, we replaced the existing mapping functions with fully connected networks, which led to a slight performance decline. This suggests that more complex network architectures are capable of learning improved mapping transformations.
@@ 52.5 $\downarrow$
Baseline Pre-train Fine-tune Avg
Exo Ego Exo Ego
VideoLLaMA2 38.9
@@ 49.1 $\uparrow$
Table 9: Effect comparison across different datasets. “Exo” refers to the partial exocentric data released by VideoLLaMA2 [11], while “Ego” encompasses both the EgoClip [49] dataset and the instruction-tuning data we constructed.
@@
Training Data. To assess the influence of data from different domains on the performance of MLLMs, we conducted experiments using VideoLLaMA2 [11] as a baseline, as summarized in Table 9. The second result indicates a significant improvement over the initial baseline, demonstrating that our egocentric instruction-tuning data EgoIT enhances the understanding of egocentric videos. In the third row of results, we replaced the pre-training dataset with EgoClip [49], which is aligned with the egocentric domain. The improved performance suggests that pre-training on data closely related to the downstream tasks leads to better domain knowledge. The comparison between results 3 and 4 is particularly thought-provoking, suggesting that domain-specific instruction data offers more substantial benefits than generalization, consistent with the findings in [47]. This underscores the considerable gap between egocentric and exocentric data.
@@
Figure 5: Accuracy comparison across different training stages.
Figure 6: Two qualitative comparisons from EgoBench.
5.4 Qualitative Comparison
To visually compare the performance of different MLLMs, we presented several qualitative examples in Figure 6. In Figure 6(a), both baselines struggle with accurately interpreting the video content. VideoLLaMA2 erroneously assumes the camera-wearer is mixing paint with a brush, while VideoChat2 mistakenly believes the brush is being cleaned. In contrast, our Exo2Ego method correctly identifies that the camera-wearer is diluting the color on the brush, showcasing its superior understanding and inference capabilities for visual details. In the second example from EgoPlan dataset, which evaluates planning abilities, our model demonstrates a clear understanding of the temporal relationships between actions. Unlike the two baselines, which provide incorrect predictions, our model accurately identifies the correct sequence of actions required to achieve the goal. For further qualitative analyses, please refer to Supp. D.3.
6 Conclusion
In this paper, we aim to improve egocentric video understanding by transferring extensive exocentric knowledge embedded in MLLMs at low cost. To achieve this, we first construct a diverse clip-text dataset, Ego-ExoClip, with synchronized ego-exo perspectives. We then design a progressive three-stage training pipeline to effectively learn the mappings between the two domains. To enhance instruction-following abilities for downstream tasks, we introduce the egocentric instruction-tuning data EgoIT. Moreover, we propose EgoBench, a benchmark designed to comprehensively evaluate the embodied cognitive abilities of existing MLLMs. Extensive experiments have validated the superiority of our Exo2Ego framework for egocentric tasks.
[…]
Supplementary Material
  • Section A: Ego-ExoClip Details
    • Section A.1: Dataset Comparison
    • Section A.2: Clip Analysis
    • Section A.3: Narration Analysis
    • Section A.4: Others
  • Section B: EgoIT Details
  • Section C: EgoBench Details
    • Section C.1: Data Collection
    • Section C.2: Example Presentation
  • Section D: More Experiments
    • Section D.1: Hyperparameter Settings
    • Section D.2: Comparison Results
    • Section D.3: Qualitative Analysis
  • Section E: Limitations
A Ego-ExoClip Details
A.1 Dataset Comparison
In Table 10, we compare our proposed Ego-ExoClip dataset with existing egocentric-exocentric datasets, including EgoExoLearn [29], Ego-Exo4D [25], and EgoExo-Fitness [47], all published in 2024. Ego-ExoClip preserves the strengths of Ego-exo4D, such as diverse activity scenarios, richly synchronized exocentric perspectives, and a large-scale dataset. Moreover, compared to Ego-Exo4D, Ego-ExoClip extends timestamp-level narrations to the clip level, enabling a broader range of pre-training tasks. Therefore, Ego-ExoClip is the largest and most diverse clip-text dataset currently available, offering synchronized egocentric-exocentric perspectives.
@@ A.1 Dataset Comparison
CharadesEgo [75] 2018 Indoor Daily Paired 1 34
LEMMA [30] 2020 Kitchen, Living Paired 1 10
HOMAGE [66] 2021 Indoor Daily Paired 2-5 25
Assembly101 [71] 2022 Desk Paired 8 42
EgoExoLearn [29] 2024 Kitchen, Lab Unpaired - 120
Ego-Exo4D [25] 2024 Diverse Paired 4-5 221
EgoExo-Fitness [47] 2024 Fitness Paired 3 32
@@ A.4 Others
Geographic Diversity. Figure 9 shows the regional distribution of the videos in Ego-ExoClip, which includes a total of 2,925 groups. The dataset covers contributions from 12 institutions across 6 different countries: India, Singapore, Japan, USA, Canada, and Colombia. This broad coverage allows our pre-training dataset to effectively inherit the rich geographic and participant diversity of Ego-Exo4D. For more detailed information, please refer to the original Ego-Exo4D paper [25].
@@ B EgoIT Details
C EgoBench Details
C.1 Data Collection
C.1.1 EgoSchema
EgoSchema [58] is a comprehensive dataset and benchmark designed to evaluate the ability of modern vision and language models to understand extended video content. Based on the Ego4D [24] dataset, EgoSchema contains over 5,000 carefully curated multiple-choice questions and answers derived from more than 250 hours of real-life video footage capturing a wide range of natural human activities and behaviors. Each question requires selecting the correct answer from five options after viewing a three-minute video segment, thus challenging models to perform detailed video-based reasoning.
@@ C.1.1 EgoSchema
C.1.2 QAEgo4D
QAEgo4D [6] is a video question-answering dataset for episodic memory tasks, designed to evaluate video comprehension and reasoning capabilities in modern vision-language systems. Sourced from Ego4D [24], it features human-annotated question-answer pairs, where each sample contains a video, a question, an answer, and the specific temporal window where the answer can be found. The dataset focuses on egocentric videos depicting natural human activities, with an average video length of 310.0 seconds in testing data. By using overlapping text descriptions for consistency, QAEgo4D serves as a valuable resource for advancing research in episodic memory and video question answering.
We use this dataset to support the episodic memory task. For the closed-set QA, we use the testing data processed by Llama2-13B [^7] https://huggingface.co/meta-llama/Llama-2-13b. from [18], while for the open-set QA, we use the original testing data from [6].
Task Source Example
@@ C.1.2 QAEgo4D
Table 12: Examples of tasks from the proposed EgoBench benchmark.
C.1.3 EgoTaskQA
@@
EgoTaskQA [31] is a large-scale egocentric video question-answering dataset designed to evaluate models’ understanding of goal-oriented human tasks. It features 40K balanced question-answer pairs derived from 2K videos in the LEMMA dataset [30], focusing on aspects such as action effects, intent, multi-agent collaboration, and object interactions. The dataset emphasizes reasoning types including spatial-temporal understanding, causal dependencies, and task planning, supported by 30K annotated state transitions. It includes a variety of question formats, such as binary and open-ended queries, to ensure a balanced and unbiased evaluation. EgoTaskQA is uniquely structured into two distinct subsets: “direct” subset contains randomly sampled questions, and “indirect” subset requires multi-step reasoning to derive the correct answer.
@@
Due to the subtlety and ambiguity of many questions, open-domain QA [92] in a zero-shot setting is difficult to accurately evaluate. Consequently, we structure it as a multiple-choice QA task. Specifically, we retain the original question from the testing set and select the four most similar incorrect answers from the answer set based on SentenceBERT [69] similarity to the ground truth. These four, along with the correct answer, form a set of five candidates.
- Metric Method
VideoLLaMA2 [11] VideoChat2 [43] Qwen2-VL [81] Exo2Ego (Ours)
ViEncoder - CLIP-L-336 InternVideo - CLIP-L-336
LLM - Mistral-7B Mistral-7B Qwen2-7B Mistral-7B
EgoSchema Acc. 51.7 54.4 59.0 61.3 $\uparrow$ 2.3
Acc. (Closed) 48.5 45.4 52.4 62.1 $\uparrow$ 9.7
QAEgo4D Acc./Score (Open) 25.3/2.6 21.9/1.7 32.6/2.9 28.3/2.7
EgoTaskQA Acc. (Direct) 42.1 42.3 51.9 44.7
Acc. (Indirect) 42.6 47.4 55.2 50.3
Charades-Ego mAP 68.1 57.0 44.2 70.9 $\uparrow$ 2.8
EK-100 MIR mAP 47.8 42.6 44.6 49.7 $\uparrow$ 1.9
nDCG 60.6 - - 63.6 $\uparrow$ 3.0
EgoPlan Val Acc. 29.3 26.1 29.4 42.7 $\uparrow$ 3.3
VLN-QA Acc. 37.5 35.0 36.5 44.5 $\uparrow$ 7.0
Acc. (Inter) 82.1 85.1 90.1 88.4
EgoMCQ Acc. (Intra) 32.0 37.6 38.6 41.2 $\uparrow$ 2.6
Table 13: Performance comparison with more SOTA methods at a uniform 16-frame input setting. The QAEgo4D dataset is categorized into “Closed” and “Open” question answering, while the EgoTaskQA dataset includes “Direct” and “Indirect” subsets. In EgoMCQ, “Inter” refers to cases where the candidates and input clips come from different videos, while “Intra” indicates that they are from the same video, presenting a more challenging scenario. The best results are highlighted in bold.
Config Initial Stage 1&2 Stage 3
Global Batch size 512 256 64
Local Batch size 8 8 8
Learning rate 1e-3 1e-4 2e-5
Warmup ratio 0.1 0.03 0.03
Total epoch 5 2 1
Table 14: Hyperparameter settings for different training stages.
C.1.4 Charades-Ego
@@
Charades-Ego [75] is a large-scale dataset comprising 7,860 videos that capture everyday indoor activities from both first- and third-person perspectives. It contains 68.8 hours of video with 68,536 temporal annotations across 157 action classes. Derived from the original Charades dataset, Charades-Ego bridges first- and third-person video understanding by offering paired videos recorded simultaneously from both perspectives. Featuring 112 actors performing a diverse range of actions, this dataset supports tasks such as video classification, localization, captioning, and cross-modal learning, making it a valuable resource for studying human activity from different viewpoints.
@@
For model evaluation, we find that a multiple-choice QA format is more reliable and stable compared to direct answer generation. Accordingly, we create a comprehensive set of all actions and select the four most similar incorrect actions to each action label using SentenceBERT [69] similarity. These four actions, combined with the correct answer, form a set of five candidate options for each sample. This approach is more effective than generating candidates with the LLMs, as it ensures that the options are realistic and require visual details for accurate differentiation.
EgoBench Metric Initial Stage 1 Stage 2 Stage 3

### v2

@@ Introduction
In the final stage, we focus on enhancing the instruction-following capabilities of MLLMs to address diverse egocentric downstream tasks. To achieve this, we employ the instruction-tuning data EgoIT. By concatenating the representations of the egocentric sample $x$ and its corresponding mapped exocentric sample $F(x)$, we create a comprehensive semantic input for the LLM. To ensure the LLM aligns its responses more effectively with task instructions, we apply Low-Rank Adaptation (LoRA) to the frozen LLM, tuning it alongside the visual encoder and the mapping function $F$ using VTG loss.
Experiment
Implementation Details
@@
Datasets. The datasets employed in each stage of training are summarized in Table 1, showcasing a substantial amount of vision-text data across various types. After training, we evaluated our model on multiple benchmarks, which includes eight tasks with multiple configurations: EgoSchema (mangalam2023egoschema) for reasoning, QAEgo4D (barmann2022did) for episodic memory, Charades-Ego (sigurdsson2018charades) for action recognition, EPIC-KITCHENS-100 (damen2022rescaling) for multi-instance retrieval, VLN-QA (krantz2020beyond) for egocentric navigation, EgoPlan (chen2023egoplan) for planning, EgoMCQ (lin2022egocentric) for alignment, and EgoTaskQA (jia2022egotaskqa) for hybrid tasks.
@@
Model. Our model is built upon the VideoLLaMA2 (cheng2024videollama) baseline. We consistently utilized the CLIP-Large-336 as the visual encoders and Mistral-7B-Instruct as the LLM. The mapping functions $F$ and $G$ consist of nine ResNet blocks between the down-sampling and up-sampling operations. In Stage 3, we incorporated LoRA into the LLM, with a rank set to 128, an alpha value of 256, and a dropout rate of 0.1. All experiments are conducted using the PyTorch framework on 16 A800 GPUs.
Baselines. We selected three categories of methods as comparative baselines, including specific approaches tailored for the egocentric domain (e.g., GroundVQA (di2024grounded) and LaViLa++ (xu2025egocentric)), closed-source MLLMs (e.g., GPT-4o), and open-source MLLMs (e.g., VideoLLaMA2 (cheng2024videollama) and Qwen2.5-VL (bai2025qwen2)). Except for LLaMA-VID (li2023llama) and Video-LLaVA (lin2023video), which use fixed input settings of 1 fps and 8 frames respectively, all other models adopt a 16-frame configuration to ensure a fair comparison.
Quantitative Evaluation
Tables 2 presents the experimental results with different superior methods across various tasks and settings, which can be analyzed as follows:
  • The experimental results on the reasoning and episodic memory tasks (i.e., the first three benchmarks) indicate that, in a zero-shot setting, MLLMs achieve performance comparable to specialized egocentric methods, highlighting their strong generalization capability across data domains. Notably, our Exo2Ego method significantly surpasses all open-source MLLMs and egocentric-specific methods under nearly all settings across the three egocentric evaluation benchmarks.
  • The results on the fourth and fifth benchmarks present zero-shot performace on action recognition (Charades-Ego) and multi-instance retrieval (EPIC-KITCHENS-100). Our Exo2Ego method demonstrates consistent improvements over all baselines across all metrics for both tasks, validating its effectiveness in understanding the details of hand-object interactions.
  • Planning and navigation are critical tasks in the domain of egocentric understanding. The sixth and seventh columns summarize our method’s performance against existing baselines. The results reveal that open-source MLLMs exhibit accuracy rates concentrated between 26.0% and 32.5% on EgoPlan, highlighting their limitations for planning. In contrast, our method significantly outperforms all competitors, including GPT-4o (5.9% and 10.5% absolute gains), demonstrating its strong capability to effectively tackle diverse egocentric tasks.
  • Video-text matching, a fundamental alignment task, is also crucial for comprehensive evaluation. As shown in the final column, our Exo2Ego approach consistently achieves the best performance across both settings, further demonstrating its superiority in generating accurate visual descriptions.
Ablation Study
In this section, we conducted a comprehensive analysis of how our model architecture, parameter update, training data, and training strategy impact overall performance.
Model Architecture. As summarized in Table 3, we investigated the impact of the mapping functions, LoRA, and loss functions on model performance. Comparing results 1 and 2, we observed that the introduction of LoRA significantly improves instruction-following capabilities, leading to enhanced performance across various tasks. This highlights LoRA’s effectiveness in refining the model’s ability to follow complex instructions. The third experiment utilizes only forward cycle consistency, resulting in a performance decline and highlighting the importance of the full bidirectional loss design. In the comparison between results 1 and 4, we found that the mapping $G$ and CCL loss play a critical role in establishing an effective mapping between the egocentric and exocentric domains, thereby improving performance on downstream egocentric tasks. In the fifth experiment, we removed the KL loss, which incorporates guidance from exocentric data. The performance gap between this result and result 1 reinforces the importance of exocentric knowledge in enhancing egocentric video understanding. In the sixth experiment, we replaced the existing mapping functions with fully connected networks, which led to a slight performance decline. This suggests that more complex network architectures are capable of learning improved mapping transformations.
@@
Table 3: Performance comparison of different model architecture. The LLM is fixed at Mistral-7B-Instruct-v0.2, “✰” denotes replacing the existing modules with fully connected networks, and “➩” represents removing backward cycle consistency, using only the forward direction. “Avg” is the average performance across all metrics in Figure 1.
@@ 52.5 $\downarrow$
EgoClip Ego-ExoClip EgoIT Avg
38.9
@@ 49.7 $\uparrow$
55.6 (Ours)
Table 5: Effect comparison across different datasets. The results in the first four rows are from VideoLLaMA2, using the same backbone as our method.
@@
Training Data. To assess the influence of training data on the model performance, we conducted experiments using VideoLLaMA2 (cheng2024videollama) as a baseline, as summarized in Table 5. Notably, VideoLLaMA2 adopts the same backbone architecture as our framework, including visual encoder and LLM. In the second row of results, we added the EgoClip dataset, which is aligned with the egocentric domain. The improved performance suggests that training on data closely related to the downstream tasks leads to better domain knowledge. Compared to result 2, result 3 incorporates the Ego-ExoClip data and achieves further performance improvements. This validates the effectiveness of our constructed dataset. The fourth result indicates a significant improvement over the third result, demonstrating that our egocentric instruction-tuning data EgoIT enhances adaptation to downstream egocentric tasks. The performance gap between the last two rows indicates that, even when using the same datasets, our framework still exhibits a significant advantage. This demonstrates the superiority of our dual-encoder architecture and transfer-based training strategy.
@@
Figure 5: Accuracy comparison across different stages.
Figure 6: Two examples from different benchmarks.
Qualitative Comparison
To visually compare the performance of different MLLMs, we presented several qualitative examples in Figure 6. In Figure 6(a), both baselines struggle with accurately interpreting the video content. VideoLLaMA2 erroneously assumes the camera-wearer is mixing paint with a brush, while VideoChat2 mistakenly believes the brush is being cleaned. In contrast, our Exo2Ego method correctly identifies that the camera-wearer is diluting the color on the brush, showcasing its superior understanding and inference capabilities for visual details. In the second example from EgoPlan dataset, which evaluates planning abilities, our model demonstrates a clear understanding of the temporal relationships between actions. Unlike the two baselines, which provide incorrect predictions, our model accurately identifies the correct sequence of actions required to achieve the goal.
Conclusion
In this paper, we propose Exo2Ego, a unified framework that significantly advances egocentric video understanding by cost-effectively transferring rich exocentric knowledge from MLLMs. To enable robust cross-view representation learning, we introduce Ego-ExoClip, a large-scale dataset comprising 1.1 million synchronized egocentric-exocentric clip-text pairs. Additionally, we curate EgoIT, an egocentric instruction-tuning dataset designed to further enhance instruction-following capabilities. By leveraging a novel exocentric-to-egocentric migration strategy and a progressive three-stage training pipeline, Exo2Ego effectively bridges the domain gap and promotes egocentric self-consistency. Extensive experiments on multiple benchmarks demonstrate that Exo2Ego consistently achieves substantial performance improvements.
— Supplementary Material —
Ego-ExoClip Details
Dataset Comparison
In Table 6, we compare our proposed Ego-ExoClip dataset with existing egocentric-exocentric datasets, including EgoExoLearn (huang2024egoexolearn), Ego-Exo4D (grauman2024ego), and EgoExo-Fitness (li2024egoexo), all published in 2024. Ego-ExoClip preserves the strengths of Ego-exo4D, such as diverse activity scenarios, richly synchronized exocentric perspectives, and a large-scale dataset. Moreover, compared to Ego-Exo4D, Ego-ExoClip extends timestamp-level narrations to the clip level, enabling a broader range of pre-training tasks. Therefore, Ego-ExoClip is the largest and most diverse clip-text dataset currently available, offering synchronized egocentric-exocentric perspectives.
@@ Conclusion
CharadesEgo 2018 Indoor Daily Paired 1 34
LEMMA 2020 Kitchen, Living Paired 1 10
HOMAGE 2021 Indoor Daily Paired 2-5 25
Assembly101 2022 Desk Paired 8 42
EgoExoLearn 2024 Kitchen, Lab Unpaired - 120
Ego-Exo4D 2024 Diverse Paired 4-5 221
EgoExo-Fitness 2024 Fitness Paired 3 32
@@
Geographic Diversity. Figure 9 shows the regional distribution of the videos in Ego-ExoClip, which includes a total of 2,925 groups. The dataset covers contributions from 12 institutions across 6 different countries: India, Singapore, Japan, USA, Canada, and Colombia. This broad coverage allows our pre-training dataset to effectively inherit the rich geographic and participant diversity of Ego-Exo4D. For more detailed information, please refer to the original Ego-Exo4D paper (grauman2024ego).
@@
Benchmark Details
Evaluation Dataset
EgoSchema
EgoSchema (mangalam2023egoschema) is a comprehensive dataset and benchmark designed to evaluate the ability of modern vision and language models to understand extended video content. Based on the Ego4D (grauman2022ego4d) dataset, EgoSchema contains over 5,000 carefully curated multiple-choice questions and answers derived from more than 250 hours of real-life video footage capturing a wide range of natural human activities and behaviors. Each question requires selecting the correct answer from five options after viewing a three-minute video segment, thus challenging models to perform detailed video-based reasoning.
@@
QAEgo4D
QAEgo4D (barmann2022did) is a video question-answering dataset for episodic memory tasks, designed to evaluate video comprehension and reasoning capabilities in modern vision-language systems. Sourced from Ego4D (grauman2022ego4d), it features human-annotated question-answer pairs, where each sample contains a video, a question, an answer, and the specific temporal window where the answer can be found. The dataset focuses on egocentric videos depicting natural human activities, with an average video length of 310.0 seconds in testing data. By using overlapping text descriptions for consistency, QAEgo4D serves as a valuable resource for advancing research in episodic memory and video question answering.
We use this dataset to support the episodic memory task. For the closed-set QA, we use the testing data processed by Llama2-13B [^3] https://huggingface.co/meta-llama/Llama-2-13b. from (di2024grounded), while for the open-set QA, we use the original testing data from (barmann2022did).
Task Benchmark Example
@@
Table 8: Examples of tasks from different benchmarks.
EgoTaskQA
@@
EgoTaskQA (jia2022egotaskqa) is a large-scale egocentric video question-answering dataset designed to evaluate models’ understanding of goal-oriented human tasks. It features 40K balanced question-answer pairs derived from 2K videos in the LEMMA dataset (jia2020lemma), focusing on aspects such as action effects, intent, multi-agent collaboration, and object interactions. The dataset emphasizes reasoning types including spatial-temporal understanding, causal dependencies, and task planning, supported by 30K annotated state transitions. It includes a variety of question formats, such as binary and open-ended queries, to ensure a balanced and unbiased evaluation. EgoTaskQA is uniquely structured into two distinct subsets: “direct” subset contains randomly sampled questions, and “indirect” subset requires multi-step reasoning to derive the correct answer.
@@
Due to the subtlety and ambiguity of many questions, open-domain QA in a zero-shot setting is difficult to accurately evaluate. Consequently, we structure it as a multiple-choice QA task. Specifically, we retain the original question from the testing set and select the four most similar incorrect answers from the answer set based on SentenceBERT (reimers2019sentence) similarity to the ground truth. These four, along with the correct answer, form a set of five candidates.
Charades-Ego
@@
Charades-Ego (sigurdsson2018charades) is a large-scale dataset comprising 7,860 videos that capture everyday indoor activities from both first- and third-person perspectives. It contains 68.8 hours of video with 68,536 temporal annotations across 157 action classes. Derived from the original Charades dataset, Charades-Ego bridges first- and third-person video understanding by offering paired videos recorded simultaneously from both perspectives. Featuring 112 actors performing a diverse range of actions, this dataset supports tasks such as video classification, localization, captioning, and cross-modal learning, making it a valuable resource for studying human activity from different viewpoints.
@@
For model evaluation, we find that a multiple-choice QA format is more reliable and stable compared to direct answer generation. Accordingly, we create a comprehensive set of all actions and select the four most similar incorrect actions to each action label using SentenceBERT (reimers2019sentence) similarity. These four actions, combined with the correct answer, form a set of five candidate options for each sample. This approach is more effective than generating candidates with the LLMs, as it ensures that the options are realistic and require visual details for accurate differentiation.
Benchmark Metric Initial Stage 1 Stage 2 Stage 3